1. The equity beta of WealthCorp Ltd is 1.10, the market risk premium is expected to be 9%, and the yield on government bonds is currently 12%. WealthCorp Ltd issued bonds (R100 par value) that are currently trading at R80 and have a 5% coupon rate. The corporate tax rate is currently 25% and the maturity date of the bonds is in 10 years. Using the CAPM, calculate the cost of equity and the market’s overall expected rate of return (Rm).


1
Expert's answer
2021-11-28T17:50:36-0500

"CAMP=0.12+1.1*0.09=0.219=29.9\%" cost of equity

"0.12+0.09=0.21" expected rate of return (Rm).